OpenStudy (anonymous):

Solve the system by graphing. x-2y=8 x-2y=4

OpenStudy (anonymous):

y=-x-4 y=-x-2 Undefined. No Solution. The lines are parallel,thus have the same slope.

OpenStudy (anonymous):

x=8+2y (8+2y)-2y=4 8=4 NNO solution

OpenStudy (anonymous):

How do you graph it then when it's undefined?

OpenStudy (anonymous):

like a regular graph.

OpenStudy (anonymous):

graph y=-x-4 then graph the other one and you will see that they are parallel